OpenStudy (anonymous):

Create a table of values for the equation y+2x=5 and graph the line. HELP

OpenStudy (lucaz):

pick some numbers for x and find the calculate the corresponding y

OpenStudy (anonymous):

What do you mean?

OpenStudy (lucaz):

the equation is y+2x=5, let's pick the number 0 y+2(0)=5 y = ?

OpenStudy (anonymous):

y=3 ?

OpenStudy (lucaz):

what is 2*0 ?

OpenStudy (anonymous):

2

OpenStudy (lucaz):

any number times zero is always zero, 2(0)=0

OpenStudy (anonymous):

Then what would y be?

OpenStudy (lucaz):

y+0=5 y=5

OpenStudy (lucaz):

this gives the point (0,5), now try with x = 1

OpenStudy (lucaz):

y+2(1)=5

OpenStudy (anonymous):

y+2=5?

OpenStudy (lucaz):

yes, now to find the value of y you subtract 2 from both sides of the equation

OpenStudy (lucaz):

y+2-2=5-2 y=?

OpenStudy (anonymous):

y=3?

OpenStudy (lucaz):

correct, this is the point (1,3)

OpenStudy (anonymous):

Thank you

OpenStudy (lucaz):

no problem